# Coordinate versus momentum cuts and effects of collective flow on critical fluctuations

### v1

@@ I Introduction
Identifying the existence and location of the QCD critical point (CP) at finite baryon density is one of the main goals of the beam energy scans performed with relativistic heavy-ion collisions Bzdak et al. (2020). Event-by-event fluctuations of the proton number are the primary observable here Stephanov et al. (1999); Hatta and Stephanov (2003). In particular, proton number cumulants are expected to show a non-monotonic collision energy dependence if the QCD critical point exists and heavy-ion collisions are sensitive to it in a narrow collision energy range Stephanov (2009). Experimental measurements performed by the STAR Collaboration within phase I of the RHIC beam energy scan show indications for a non-monotonic collision energy dependence of the proton kurtosis $\kappa\sigma^{2}=\kappa_{4}/\kappa_{2}$ Adam et al. (2021), although the experimental error bars are still too large to draw firm conclusions. On the other hand, 2nd order cumulants of protons were measured with much larger precision Abdallah et al. (2021), and show indications for an excess of the proton number scaled variance at $\sqrt{s_{\rm NN}}\lesssim 20$ GeV relative to baseline expectations due to baryon number conservation and repulsive baryon hard-core (see Vovchenko (2023a) for a recent overview). Interestingly, measurements at even lower energies, $\sqrt{s_{\rm NN}}=2.4$ GeV by HADES Adamczewski-Musch et al. (2020) and $\sqrt{s_{\rm NN}}=3$ GeV by STAR Abdallah et al. (2022) also show indications for the large variance of the proton number distribution, although these measurements are affected by large volume fluctuation effects unrelated to the CP. Therefore, the effort to locate the CP with heavy-ion collisions is now mainly focused on collision energies of $\sqrt{s_{\rm NN}}\simeq 2.4-20$ GeV, with future experimental data coming from RHIC BES-II and fixed target programs, as well as the CBM experiment at FAIR Ablyazimov et al. (2017). In addition, several recent effective QCD approaches Ratti et al. (2006); Ding et al. (2016); Fu et al. (2020); Gunkel and Fischer (2021); Hippert et al. (2023); Basar (2023); Goswami et al. (2024) constrained by lattice QCD simulations at $\mu_{B}=0$ place the CP into a $T-\mu_{B}$ range probed by intermediate energy heavy-ion collisions Vovchenko (2023a).
@@ II.2 Molecular dynamics
MD simulations proceed by numerically integrating Newton’s equations of motion. The simulations are performed using the Velocity-Verlet integration method for the system of $N$ particles with periodic boundary conditions in the minimum-image convention form [^1] One can see details of method in Allen and Tildesley (2017) and find the simulation setup source in Vovchenko (2023b). In the previous work Kuznietsov et al. (2022), we used the simulations to study the behavior of particle number fluctuations along the supercritical isotherm $\tilde{T}=1.06\tilde{T}_{c}$. This was achieved by performing the simulations for a long period of time at each value of particle number densities and computing the moments of particle number distribution as time average.
In the present work, we explore the same conditions of temperature and density as in Ref. Kuznietsov et al. (2022) and use the same GPU-accelerated MD simulation code from Vovchenko (2023b). We refer to Sec. III of Ref. Kuznietsov et al. (2022) for the details of MD simulation framework. The key difference to Ref. Kuznietsov et al. (2022) is that here we calculate the observables as ensemble averages, namely, by performing a large number of MD simulations at each density, each simulation initialized with random initial conditions. In this way we are able to compare ensemble averaging with time averaging in Ref. Kuznietsov et al. (2022) and study equilibration dynamics at different conditions of particle number density. Our simulations here are performed for $N=400$, which approximately corresponds to the total number of baryons in central collisions of heavy ions when the production of baryon-antibaryon pairs is negligible.
@@ II.3.3 MD simulation
Each event is propagated from the initial time $\tilde{t}=0$ to $\tilde{t}=100$ by solving the equations of motion with the GPU-accelerated MD solver Vovchenko (2023b). We use a time step size of $\Delta\tilde{t}=0.004$ for $\tilde{n}=0.32\tilde{n}_{\rm c}$ and $0.95\tilde{n}_{\rm c}$, and a smaller value of $\Delta\tilde{t}=0.002$ for $\tilde{n}=1.9\tilde{n}_{\rm c}$. These values were found to be sufficient to maintain the numerical stability and accuracy of the simulations, which we verified by monitoring the conservation of energy $\tilde{U}$ throughout the simulation.
@@ IV.1 Incorporating longitudinal flow
The collective component of the rapidity is a function of coordinate. Due to the ultrarelativsitic nature of the motion in the longitudinal direction, it is common to work in Bjorken variable, where instead of the Euclidean time $t$ and the longitudinal coordinate $z$ one uses Milne coordinates, the longitudinal proper time $\tau=\sqrt{t^{2}-z^{2}}$ and the space-time rapidity, $\eta_{s}=\frac{1}{2}\ln\frac{t-z}{t+z}$, are used instead. In the Bjorken-like longitudinal flow picture, collective component of the rapidity coincides with the space-time rapidity $y_{\rm coll}=\eta_{s}$. To make a connection between the LJ longitudinal coordinate $\tilde{z}^{\rm LJ}$ and $\eta_{s}$ at a given collision energy $\sqrt{s_{\rm NN}}$, we make a linear map between the interval $\tilde{z}^{\rm LJ}\in[-\tilde{L}/2,\tilde{L}/2]$ and the space-time rapidity extent $\eta_{s}\in[-y_{\rm cm}^{\rm beam},y_{\rm cm}^{\rm beam}]$, where
@@ IV.3 Fluctuations at fixed $y_{\rm cut}$
Remarkably, our results indicate that the maximum CP signal would be observed at $\sqrt{s_{\rm NN}}\sim 5-7$ GeV, i.e. if the CP is accessible in heavy-ion regime, these collision energies are optimal for observing its signatures in baryon number cumulants. This ’sweet spot’ in $\sqrt{s_{\rm NN}}$ is an interplay of two effects. At lower collision energies, the CP signal is diluted due to a weak collective flow and the absence of correlations between coordinates and momenta. On the other hand, larger collision energies correspond to smaller values of $\alpha$ [^6] In the limit $\sqrt{s_{\rm NN}}\to\infty$ we have $\alpha\to 0$ and thus $\tilde{\omega}_{\alpha}\to 1$.. Thus, the finite-size effects are stronger at large $\sqrt{s_{\rm NN}}$, and these dampen the CP signal (see Fig. 3). Interestingly, a similar, broader maximum is observed for $n=0.32n_{c}$, where the fluctuations also show enhancement, albeit smaller in magnitude compared to the CP. At large density, $n=1.9n_{c}$, the fluctuations are suppressed due to repulsive interactions, this suppression decreases monotonically with collision energy.
@@
We note that baryon fluctuations are not measured in the experiment directly. Instead, we reconstruct $\omega_{B}$ from the measured $\omega_{p}$ through the unfolding method from Kitazawa and Asakawa (2012). We perform this reconstruction for a meaningful comparison of the data with our model predictions, which are performed for all baryons rather than for protons only.
The experimental data show enhancement of $\tilde{\omega}$ with respect to unity at low collision energies and suppression at large energies. The maximum value of $\tilde{\omega}_{B}\simeq 1.15$ is reached at lowest available BES energy of $\sqrt{s_{\rm NN}}=7.7$ GeV. This indicates enhancement of fluctuations, although it is still considerably lower than our equilibrium predictions for $n=0.32n_{c}$ and $n=0.95n_{c}$.
@@ IV.4 Finite-time effects
Here we address this question, focusing on fluctuations in the rapidity space for fixed $y_{\rm cut}=0.5$ and fixed energy of $\sqrt{s_{\rm NN}}=7.7$ GeV. Figure 9 depicts the time dependence of $\tilde{\omega}$. This quantity reaches the equilibrium expectation shown by the red band at $\tilde{t}\gtrsim 10$. At shorter times, however, large deviations from the equilibrium value are seen. In particular, this is the case for $\tilde{t}\sim 3-5$ relevant for heavy-ion collisions, as discussed above. Of course, given the limitations of our approach, in particular the difference between initial conditions in our simulations and those in heavy-ion collisions, these estimates can give only a qualitative picture for he possible finite-time effects.
@@
Finite-time effects have a very mild effect on the calculation at $n=1.9n_{c}$ given that the time $\tilde{t}=3$ is considerably larger than the corresponding equilibration time for fluctuations, which for $n=1.9n_{c}$ is $\tilde{\tau}_{\rm eq}\sim 1$ (Table 1).
@@ IV.5 Acceptance dependence
The experimental data are available up to $y_{\rm cut}=0.5$, shown in Fig. 10 for the reconstructed baryons. The comparison with calculations results indicates a better agreement of the $n=0.32n_{c}$ case at $y_{\rm cut}\leq 0.4$ over $n=0.95n_{c}$. Both cases predict continued growth up to $y_{\rm cut}=1$, which can be verified with BES-II measurements utilizing expanded rapidity coverage.
@@ IV.6 Limitations
It is important to emphasize the limitations of our approach when applied to measurements from heavy-ion collisions. The main caveat is that we do not simulate the full dynamics of heavy-ion collisions but perform box simulations of the subsystem of nucleons near the CP, and implement the collective flow effect on top of these simulations. It is thus assumed that our LJ molecular dynamics simulation models the behavior of (critical) fluctuations nucleons in the local rest frame, where we rely heavily on the universality of critical behavior, while collective expansion is described through a separate mechanism.
@@
Due to the above limitations, our results should mainly be viewed as qualitative expectations for the possible CP signals in fluctuations, which nevertheless do include such essential effects as exact baryon conservation, difference between coordinate and momentum space, and finite-size and finite-time effects. Our description of heavy-ion can be made more quantitative in the future by incorporating the interactions responsible for critical fluctuation dynamics into transport model framework such as SMASH Weil et al. (2016) or UrQMD Bass et al. (1998); Bleicher et al. (1999).
@@ V Discussion and summary
It should be emphasized that our present approach needs further improvements for more quantitative applications to experimental measurements. In particular, the system in our simulations evolves at a constant particle number and energy density, which is not the case for heavy-ion collisions. As mentioned before, we also neglect the production of antibaryons, which would be required for applications at $\sqrt{s_{\rm NN}}\gtrsim 20$ GeV. The modeling can also be improved by considering more realistic longitudinal density and flow profiles (as opposed to the Bjorken-like picture employed here), as well as transverse expansion and $p_{T}$ cuts. These extensions will be the subject of future studies.

### v2

@@ I Introduction
Identifying the existence and location of the QCD critical point (CP) at finite baryon density is one of the main goals of the beam energy scans performed with relativistic heavy-ion collisions Bzdak et al. (2020). Event-by-event fluctuations of the proton number are the primary observable here Stephanov et al. (1999); Hatta and Stephanov (2003). In particular, proton number cumulants are expected to show a non-monotonic collision energy dependence if the QCD critical point exists and heavy-ion collisions are sensitive to it in a narrow collision energy range Stephanov (2009). Experimental measurements performed by the STAR Collaboration within phase I of the RHIC beam energy scan show indications for a non-monotonic collision energy dependence of the proton kurtosis $\kappa\sigma^{2}=\kappa_{4}/\kappa_{2}$ Adam et al. (2021), although the experimental error bars are still too large to draw firm conclusions. On the other hand, 2nd order cumulants of protons were measured with much larger precision Abdallah et al. (2021), and show indications for an excess of the proton number scaled variance at $\sqrt{s_{\rm NN}}\lesssim 20$ GeV relative to baseline expectations due to baryon number conservation and repulsive baryon hard-core (see Vovchenko, Volodymyr (2024) for a recent overview). Interestingly, measurements at even lower energies, $\sqrt{s_{\rm NN}}=2.4$ GeV by HADES Adamczewski-Musch et al. (2020) and $\sqrt{s_{\rm NN}}=3$ GeV by STAR Abdallah et al. (2022) also show indications for the large variance of the proton number distribution, although these measurements are affected by large volume fluctuation effects unrelated to the CP. Therefore, the effort to locate the CP with heavy-ion collisions is now mainly focused on collision energies of $\sqrt{s_{\rm NN}}\simeq 2.4-20$ GeV, with future experimental data coming from RHIC BES-II and fixed target programs, as well as the CBM experiment at FAIR Ablyazimov et al. (2017). In addition, several recent effective QCD approaches Ratti et al. (2006); Ding et al. (2016); Fu et al. (2020); Gunkel and Fischer (2021); Hippert et al. (2023); Başar (2024); Goswami, Jishnu et al. (2024) constrained by lattice QCD simulations at $\mu_{B}=0$ place the CP into a $T-\mu_{B}$ range probed by intermediate energy heavy-ion collisions Vovchenko, Volodymyr (2024).
@@ II.2 Molecular dynamics
MD simulations proceed by numerically integrating Newton’s equations of motion. The simulations are performed using the Velocity-Verlet integration method for the system of $N$ particles with periodic boundary conditions in the minimum-image convention form [^1] One can see details of method in Allen and Tildesley (2017) and find the simulation setup source in Vovchenko (2023). In the previous work Kuznietsov et al. (2022), we used the simulations to study the behavior of particle number fluctuations along the supercritical isotherm $\tilde{T}=1.06\tilde{T}_{c}$. This was achieved by performing the simulations for a long period of time at each value of particle number densities and computing the moments of particle number distribution as time average.
In the present work, we explore the same conditions of temperature and density as in Ref. Kuznietsov et al. (2022) and use the same GPU-accelerated MD simulation code from Vovchenko (2023). We refer to Sec. III of Ref. Kuznietsov et al. (2022) for the details of MD simulation framework. The key difference to Ref. Kuznietsov et al. (2022) is that here we calculate the observables as ensemble averages, namely, by performing a large number of MD simulations at each density, each simulation initialized with random initial conditions. In this way we are able to compare ensemble averaging with time averaging in Ref. Kuznietsov et al. (2022) and study equilibration dynamics at different conditions of particle number density. Our simulations here are performed for $N=400$, which approximately corresponds to the total number of baryons in central collisions of heavy ions when the production of baryon-antibaryon pairs is negligible.
@@ II.3.3 MD simulation
Each event is propagated from the initial time $\tilde{t}=0$ to $\tilde{t}=100$ by solving the equations of motion with the GPU-accelerated MD solver Vovchenko (2023). We use a time step size of $\Delta\tilde{t}=0.004$ for $\tilde{n}=0.32\tilde{n}_{\rm c}$ and $0.95\tilde{n}_{\rm c}$, and a smaller value of $\Delta\tilde{t}=0.002$ for $\tilde{n}=1.9\tilde{n}_{\rm c}$. These values were found to be sufficient to maintain the numerical stability and accuracy of the simulations, which we verified by monitoring the conservation of energy $\tilde{U}$ throughout the simulation.
@@ IV.1 Incorporating longitudinal flow
The collective component of the rapidity is a function of coordinate. Due to the ultrarelativsitic nature of the motion in the longitudinal direction, it is common to work in Bjorken variable, where instead of the Minkowski time $t$ and the longitudinal coordinate $z$ one uses Milne coordinates, the longitudinal proper time $\tau=\sqrt{t^{2}-z^{2}}$ and the space-time rapidity, $\eta_{s}=\frac{1}{2}\ln\frac{t-z}{t+z}$, are used instead. In the Bjorken-like longitudinal flow picture, collective component of the rapidity coincides with the space-time rapidity $y_{\rm coll}=\eta_{s}$. To make a connection between the LJ longitudinal coordinate $\tilde{z}^{\rm LJ}$ and $\eta_{s}$ at a given collision energy $\sqrt{s_{\rm NN}}$, we make a linear map between the interval $\tilde{z}^{\rm LJ}\in[-\tilde{L}/2,\tilde{L}/2]$ and the space-time rapidity extent $\eta_{s}\in[-y_{\rm cm}^{\rm beam},y_{\rm cm}^{\rm beam}]$, where
@@
The procedure described above is the simplest one for implementing longitudinal flow into the system, which corresponds to the transformation of a single fireball in a box into an expanding one (see Fig. 4). It relies on the Bjorken picture [the second term in Eq. (9)] as well as the absence of event-by-event fluctuations of in the longitudinal flow. As such, the description must be improved for quantitative applications, especially at lower energies from RHIC beam energy scan. In the present work, we retain the picture presented above to make a first estimate of the effect of longitudinal flow under the most favorable (and simplified) conditions possible.
@@ IV.3 Fluctuations at fixed $y_{\rm cut}$
Our calculations indicate that the maximum CP signal would be observed at $\sqrt{s_{\rm NN}}\sim 5-7$ GeV, i.e. if the CP is accessible in heavy-ion regime, these collision energies are optimal for observing its signatures in baryon number cumulants based on our description. This ’sweet spot’ in $\sqrt{s_{\rm NN}}$ is an interplay of two effects. At lower collision energies, the CP signal is diluted due to a weak collective flow and the absence of correlations between coordinates and momenta. On the other hand, larger collision energies correspond to smaller values of $\alpha$ [^6] In the limit $\sqrt{s_{\rm NN}}\to\infty$ we have $\alpha\to 0$ and thus $\tilde{\omega}_{\alpha}\to 1$.. Thus, the finite-size effects are stronger at large $\sqrt{s_{\rm NN}}$, and these dampen the CP signal (see Fig. 3). Interestingly, a similar, broader maximum is observed for $n=0.32n_{c}$, where the fluctuations also show enhancement, albeit smaller in magnitude compared to the CP. At large density, $n=1.9n_{c}$, the fluctuations are suppressed due to repulsive interactions, this suppression decreases monotonically with collision energy. Of course, given the limitations of our approach (Sec. IV.6) and the complexity of the system created in heavy-ion collisions at moderate energies, our conclusions could be modified in a more involved approach. However, we do expect the qualitative interplay between the increase of system size at fixed $y_{\rm cut}$ and the dilution of space-momentum correlations as $\sqrt{s_{\rm NN}}$ is decreased to hold.
@@
We note that baryon fluctuations are not measured directly in the experiment. Instead, we reconstruct $\omega_{B}$ from the measured $\omega_{p}$ through the unfolding method from Kitazawa and Asakawa (2012). We perform this reconstruction to ensure a meaningful correspondence between the measured quantities and those computed in our model. The experimental data show enhancement of $\tilde{\omega}$ with respect to unity at low collision energies and suppression at large energies. The maximum value of $\tilde{\omega}_{B}\simeq 1.15$ is reached at the lowest available BES energy of $\sqrt{s_{\rm NN}}=7.7$ GeV. This indicates enhancement of fluctuations, although the data are considerably closer to the baseline than our equilibrium calculations for $n=0.32n_{c}$ and $n=0.95n_{c}$. We would like to emphasize here, however, that our model is not sufficiently sophisticated to draw conclusions from quantitative comparisons with experimental data. Instead, we make comparisons with data in Fig. 8 to study qualitative behavior, as well as to estimate the possible magnitude of the CP signal in heavy-ion collisions under the most favorable conditions possible.
@@ IV.4 Finite-time effects
Here we address this question, focusing on fluctuations in the rapidity space for fixed $y_{\rm cut}=0.5$ and fixed energy of $\sqrt{s_{\rm NN}}=7.7$ GeV. Figure 9 depicts the time dependence of $\tilde{\omega}$. This quantity reaches the equilibrium expectation shown by the red band at $\tilde{t}\gtrsim 10$. At shorter times, however, large deviations from the equilibrium value are seen. In particular, this is the case for $\tilde{t}\sim 3-5$ relevant for heavy-ion collisions, as discussed above. Of course, given the limitations of our approach, in particular the difference between initial conditions in our simulations and those in heavy-ion collisions, these estimates can give only a qualitative picture of the possible finite-time effects.
@@
Finite-time effects have a very mild effect on the calculation at $n=1.9n_{c}$ given that the time $\tilde{t}=3$ is considerably larger than the corresponding equilibration time for fluctuations, which for $n=1.9n_{c}$ is $\tilde{\tau}_{\rm eq}\sim 1$ (Table 1). Qualitatively, one can draw a conclusion that a system dominated by repulsive interactions shows a suppression of fluctuations, which locally equilibrate on faster time scales than those driven by attractive interactions.
@@ IV.5 Acceptance dependence
The experimental data are available up to $y_{\rm cut}=0.5$, shown in Fig. 10 for the reconstructed baryons. The data agrees qualitatively with both the $n=0.32n_{c}$ and $n=0.95n_{c}$ cases, indicating that it may be challenging to distinguish whether the system is close to the CP or not. Quantitatively, a better agreement is seen for the $n=0.32n_{c}$ case, however, we must emphasize again here that our model is not yet suited for drawing robust quantitative conclusions. Both cases predict continued growth up to $y_{\rm cut}=1$, which can be verified with BES-II measurements utilizing expanded rapidity coverage.
@@ IV.6 Limitations
It is important to emphasize the limitations of our approach when applied to measurements from heavy-ion collisions. The main caveat is that we do not simulate the full dynamics of heavy-ion collisions but perform box simulations of the subsystem of nucleons near the CP, and implement the collective flow effect on top of these simulations. It is thus assumed that our LJ molecular dynamics simulation models the behavior of (critical) fluctuations nucleons in the local rest frame, where we rely heavily on the universality of critical behavior, while collective expansion is described through a separate mechanism. In particular, this implies that, in our approach, particles in a box reach equilibrium [^10] In the case of a finite-time calculation ($\tilde{\tau}=3$), the equilibrium may be incomplete. first and then the shift of $y^{\rm acc}$ is applied. On the other hand, local equilibrium may be maintained at best in heavy-ion collisions, while the system never reaches a global equilibrium due to the continued expansion. Furthermore, we also neglect event-by-event fluctuations of the longitudinal flow, as well as fluctuations and the inhomogeneous rapidity distribution of baryon stopping.
@@
Due to the above limitations, our results should mainly be viewed as qualitative expectations for the possible CP signals in fluctuations, which nevertheless do include such essential effects as exact baryon conservation, difference between coordinate and momentum space, and finite-size and finite-time effects.
Our implementation can be improved in different ways to make the predictions more quantitative. For instance, instead of a single box, we could consider a collection of boosted fireballs along the longitudinal axis, each described by a separate LJ system in the local rest frame. This could mimick better the local equilibrium of the expanding system in heavy-ion collisions but will have to be accompanied by an analysis of the meaningful box size, conservation laws, cross-talk between the boxes, and so on. Another possibility would be incorporating the interactions responsible for critical fluctuation dynamics into transport model framework such as SMASH Weil et al. (2016) or UrQMD Bass et al. (1998); Bleicher et al. (1999).
@@ V Discussion and summary
It should be emphasized that our present approach needs further improvements for more quantitative applications to experimental measurements. In particular, the system in our simulations evolves at a constant particle number and energy density, which is not the case for heavy-ion collisions. As mentioned before, we also neglect the production of antibaryons, which would be required for applications at $\sqrt{s_{\rm NN}}\gtrsim 20$ GeV. The modeling can also be improved by considering more realistic longitudinal density and flow profiles (as opposed to the Bjorken-like picture employed here), their event-by-event fluctuations, as well as incorporating transverse expansion and $p_{T}$ cuts. One can also implement local equilibrium relevant for heavy-ion collisions by considering a collection of boosted fireballs along the longitudinal axis instead of a single fireball. Each fireball would be described by a separate LJ system in the local rest frame and will have to be accompanied by an analysis of the meaningful box size, conservation laws, cross-talk between the boxes etc. These extensions will be the subject of future studies.